OpenStudy (anonymous):

According to data on the state of Iowa, the amount of water used for cooking per household per month is normally distributed, with a standard deviation of 10.4 liters. A survey was carried out to estimate the population mean. With a sample size of 60 households, the mean quantity of water used for cooking was found to be 62.6 liters. What can we say about the population mean? 1 There is a 95% chance that the population mean would lie between 59.91 liters and 65.29 liters. 2 There is a 95% chance that the population mean would lie between 61.11 liters and 65.09 liters. @amistre64 I think it may be answer 1.

OpenStudy (anonymous):

3 There is a 68% chance that the population mean would lie between 61.26 liters and 64.49 liters. 4 There is a 68% chance that the population mean would lie between 62.11 liters and 63.09 liters.

OpenStudy (anonymous):

Do you know how to solve it?

OpenStudy (anonymous):

I think it's number 1.

OpenStudy (amistre64):

what makes you think i know this lol, let me read thru it

OpenStudy (amistre64):

the amount of water used for cooking per household per month is normally distributed, with a standard deviation of 10.4 liters. A survey was carried out to estimate the population mean. sample size of 60, mean 62.6 liters. What can we say about the population mean? what value of z is associated with the options? 95% and 68% im thinking 2 and 1 by empirical rule for approximations

OpenStudy (amistre64):

mean +- z(10.4)/sqrt(60) gives us the confidence interval related to the z value for the confidence

OpenStudy (amistre64):

\[CI_{95}\approx 62.6\pm 2\frac{10.4}{\sqrt{60}}\] \[CI_{68}\approx 62.6\pm 1\frac{10.4}{\sqrt{60}}\]

OpenStudy (amistre64):

3 would have an upper limit of 63.94 1 would have an upper limit of 65.28 -------------------- 1 seems more in line then 3 so mee